OpenStudy (anonymous):

Help please! Boyle's law for confined gases states that if the temperature is constant, pv=c where p is pressure,V is volume, and c is a constant. At a certain instant the volume is 75 cubic inches, the pressure is 3 psi, and the pressure is decreasing at the rate of 2 psi every minute. What is the rate of change of the volume at that instant?

OpenStudy (anonymous):

This is a calculus my math professor gave the class and I never could get the answer. He worked it out but it made absolutely no sense...Can some one show me how to get the answer?

zepdrix (zepdrix):

Oh rate of change :) these are fun.

OpenStudy (anonymous):

Do you think you can solve it? Because Im clueless at this point :P

zepdrix (zepdrix):

Yah just a sec though, I don't wanna interrupt abbot if he's got this :3

OpenStudy (anonymous):

Lol okie dokie :D

OpenStudy (abb0t):

It might make it easier to write down all your known's first: KNOWN: V = 75 cubic inches P = 3 \[\frac{ dP }{ dt } = -2\] negative because it's decreasing. UNKNOWN: \[\frac{ dV }{ dt }=?\] Solve for the above by differentiating w/ respect to time. does that help?

OpenStudy (abb0t):

[Note: PV should be differentiated by using the product rule! Also, C = a constant, so it becomes 0]

OpenStudy (anonymous):

So basically the prime of P times the second plus the first times the prime of the second?

OpenStudy (abb0t):

\[P (\frac{ dV }{ dt })+v(\frac{ dP }{ dt })\]

OpenStudy (anonymous):

So 3(dV/dt) + 75(-2)

OpenStudy (anonymous):

Then I just do a little algebra and get dV/dt? correct?

OpenStudy (abb0t):

Yep. Now, use algebra to rearrange and solve for dV/dt.

OpenStudy (anonymous):

So dV/dt equals 50?

OpenStudy (anonymous):

Is that the final answer or is there more to this problem?

OpenStudy (phi):

So dV/dt equals 50? For these kinds of "physics" problems I would include the units \[ \frac{dV}{dt}= +50 \text{ in}^3/\text{min} \] the balloon is getting bigger by that amount every minute, but this rate will change as the volume gets bigger, the pressure goes lower, and the leak slows down.